OpenStudy (anonymous):

Simplify: 25a^3b^2 ------------- 45a^2b^4 Answers: 5ab^2 -------- 9 5a -------- 9b^2 0.5ab^2

OpenStudy (anonymous):

5a -------- 9b^2

OpenStudy (john_es):

\[\frac{25a^3b^2}{45a^2b^4}=\frac{5^2a^2b^2\cdot a}{5a^2b^2\cdot9b^2} =\frac{5a}{9b^2}\]